OpenStudy (anonymous):

Determine whether the system of linear equations has one and only one solution, infinitely many solutions, or no solution. 2x−3y=6 4x−6y=18

OpenStudy (anonymous):

how do i start?

OpenStudy (anonymous):

Okay 2 ways, the graphical method: which is just to arrange both equations to the subject of y. meaning 2x-3y=6 becomes y= (2/3)x -2 and so on... then typing both equations into the GC, see if the 2 lines intersect, if it doesnt that means there are no solutions. The 2nd way is just to solve the 2 equations. 2x−3y=6 => y = (2x-6)/3 4x−6y=18 => y= (4x-18)/6 Equating both eqn tgt (2x-6)/3 = (4x-18)/6 4x-12 = 4x- 18 -12= -18 Which is a contradiction, so since there are no values of x that can fit into the eqn, there is no solution

OpenStudy (anonymous):

can you just explain to me why did you put 2/3x-2?

OpenStudy (anonymous):

Oh that also means y = (2x-6)/3, i just simplified it. sry for the confusion

OpenStudy (anonymous):

thank you! i was like waiit a minute that makes no sense